OpenStudy (anonymous):

A cyclist pedals along a straight road with velocity v(t)=12t^2-48t+36 (in mph) for 0

OpenStudy (anonymous):

this is a strange way of asking solve \(12t^2-48t+36>0\)

OpenStudy (anonymous):

because if the velocity is positive, that is the same as saying she moves in the positive direction start by dividing by 12 and solve \[t^2-4t+3>0\]

OpenStudy (anonymous):

+ve 3<t<1 -ve 1<t<3

OpenStudy (anonymous):

ok so on the interval \(0\le t\le 3\) it is negative on \((1,3)\) and positive on \((0,1)\) btw there is no such interval as \(3<t<1\) it is two intervals \((-\infty, 1)\cup (3,\infty)\)

OpenStudy (anonymous):

@satellite73 can you show me how you got the answers

OpenStudy (anonymous):

\(t^2-4t+3=(t-1)(t-3)\) is zero at \(t=1,t=3\) since this is a parabola that opens up, it is negative between the zeros and positive outside of them 

OpenStudy (phi):

if you want to solve it algebraically (t-1)(t-3) > 0 when both terms are positive: (t-1)> 0 and (at the same time) t-3>0 this requires t>1 and t>3, which simplifies to t>3 or both terms are negative t-1<0 and t-3<0 t< 1 and t<3. this becomes just t<1 so positive v for t<1 or t>3